OpenStudy (anonymous):

Need Geometry help! Indicate in standard form the equation of the line passing through the given point and having the given slope. R(4, 0), m = 5

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

can you help me plug it in so i can understand it better it's still confusing

jimthompson5910 (jim_thompson5910):

m is your slope (x1,y1) is the point the line goes through

jimthompson5910 (jim_thompson5910):

m = 5 is given as the slope (4,0) is the given point, so (x1,y1) = (4,0) ----> x1 = 4 and y1 = 0

OpenStudy (anonymous):

it wants me to answer in a box where the second position is taken by an x and the 7th by a 0

OpenStudy (anonymous):

so i'm still not getting how to plug it all in

jimthompson5910 (jim_thompson5910):

does the formula y - y1 = m(x - x1) look familiar?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok all you do is replace the letters with the known and given info

jimthompson5910 (jim_thompson5910):

for instance, you know that m = 5, so you replace m with 5 y - y1 = m(x - x1) y - y1 = 5(x - x1) ... replace m with 5 (or plug in m = 5)

jimthompson5910 (jim_thompson5910):

do the same for x1 and y1

OpenStudy (anonymous):

so 0=5*4 but that doesn't make any sense

jimthompson5910 (jim_thompson5910):

y - y1 = 5(x - x1) y - y1 = 5(x - 4) ... replace x1 with 4 y - 0 = 5(x - 4) ... replace y1 with 0

jimthompson5910 (jim_thompson5910):

see how I'm getting all this?

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

your next move is to get the equation info Ax+By = C form

OpenStudy (anonymous):

how do you do that?

jimthompson5910 (jim_thompson5910):

I'll give a similar example

jimthompson5910 (jim_thompson5910):

y - 2 = 3(x-7) y - 2 = 3x-21 ... distribute y - 2+2 = 3x - 21+2 ... Add 2 to both sides. y = 3x - 19 ... Combine like terms. y - 3x = 3x-19-3x ... Subtract 3x from both sides. y - 3x = -19 ... Combine like terms. -3x + y = -19 3x - y = 19 ... multiply everything by -1 this is NOT the answer you want since it's just an example

OpenStudy (anonymous):

ok so i ended up with y-5x=-20

OpenStudy (anonymous):

or -5x=-y-20

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

y-5x=-20 or -5x + y = -20 are correct

OpenStudy (anonymous):

but that still doesn't add up because the box i have to input this into has the second position taken by an x

OpenStudy (anonymous):

so theres no room for the negative

jimthompson5910 (jim_thompson5910):

some books or teachers will want A > 0, so that means you multiply everything by -1 to get -5x + y = -20 -1*(-5x) + -1*y = -1*(-20) 5x - y = 20

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

im going to see if this works

jimthompson5910 (jim_thompson5910):

it should or there's a typo somewhere

OpenStudy (anonymous):

yes perfect thanks very much

jimthompson5910 (jim_thompson5910):

yw